OpenStudy (anonymous):

(Inverse) y=(1-x^1/2)/(1+x^1/2)

OpenStudy (anonymous):

\[x=\frac{ 1-\sqrt{y} }{ 1+\sqrt{y} }\]

OpenStudy (anonymous):

Multiply by Conject?

OpenStudy (anonymous):

\[x=\frac{ 1-2\sqrt{y}+y }{ 1-y }\]

OpenStudy (anonymous):

\[x=\frac{ -2\sqrt{y} + y }{ -y}\]

OpenStudy (anonymous):

is this the right path?

OpenStudy (anonymous):

\[x+1 = \frac{ 2\sqrt{y} }{ y }\]

OpenStudy (anonymous):

\[y(x+1)=2\sqrt{y}\]

OpenStudy (anonymous):

\[\frac{ y(x+1) }{ 2 } = \sqrt{y}\]

OpenStudy (anonymous):

Is this looking right?

OpenStudy (anonymous):

The trick to finding the inverse is to go back to the original, substitute x for y, and y for x. Your new problem would be\[x=\frac{ 1-y ^{\frac{ 1 }{ 2 }} }{ 1+x ^{\frac{ 1 }{ 2 }} }\]Now solve for y

OpenStudy (anonymous):

would it be /1+y^(1/2)?

OpenStudy (anonymous):

I'm sorry made a typo RHS denominator should be y to half power

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Would I cross multiply or is there something "simplier" to do?

OpenStudy (anonymous):

????? \[x(1+y ^{1/2}) = 1-y ^{1/2}\]

OpenStudy (anonymous):

\[x+xy ^{1/2} = 1 - y ^{1/2}\]

OpenStudy (anonymous):

It's a clever problem. You have to put the apples on one side and the oranges on the other side.

OpenStudy (anonymous):

\[y ^{1/2} +xy ^{1/2} = 1-x\]

OpenStudy (anonymous):

\[y ^{1/2} = \frac{ 1-x }{ 1+x }\]

OpenStudy (anonymous):

\[y=\frac{ (1-x)^{2} }{(1+x)^{2}}\]

OpenStudy (anonymous):

???? /Holds breath

OpenStudy (anonymous):

Yeah, that's right, good job.

OpenStudy (anonymous):

Wooohooo Thank you!